Let f(x)=2x^2-7x+5 and g(x)=x^2+4. Find f*g. Simplify your answers if possible.
f*g=

Answer by Cromlix(4381)

Hi there,
Let f(x)=2x^2-7x+5 and g(x)=x^2+4. Find f*g.
f*g =
x^2 + 4(2x^2 - 7x + 5)
2x^4 - 7x^3 + 5x^2 + 8x^2 - 28x + 20
2x^4 - 7x^3 + 13x^2 - 28x + 20
Hope this helps :-)